### v1

# From Similarity to Superiority: Channel Clustering for Time Series Forecasting

Abstract
Time series forecasting has attracted significant attention in recent decades. Previous studies have demonstrated that the Channel-Independent (CI) strategy improves forecasting performance by treating different channels individually, while it leads to poor generalization on unseen instances and ignores potentially necessary interactions between channels. Conversely, the Channel-Dependent (CD) strategy mixes all channels with even irrelevant and indiscriminate information, which, however, results in oversmoothing issues and limits forecasting accuracy. There is a lack of channel strategy that effectively balances individual channel treatment for improved forecasting performance without overlooking essential interactions between channels. Motivated by our observation of a correlation between the time series model’s performance boost against channel mixing and the intrinsic similarity on a pair of channels, we developed a novel and adaptable Channel Clustering Module (CCM). CCM dynamically groups channels characterized by intrinsic similarities and leverages cluster identity instead of channel identity, combining the best of CD and CI worlds. Extensive experiments on real-world datasets demonstrate that CCM can (1) boost the performance of CI and CD models by an average margin of $2.4\%$ and $7.2\%$ on long-term and short-term forecasting, respectively; (2) enable zero-shot forecasting with mainstream time series forecasting models; (3) uncover intrinsic time series patterns among channels and improve interpretability of complex time series models.
Machine Learning, ICML
[…]
1 Introduction
Figure 1: The pipeline of applying Channel Clustering Module (CCM) to general time series models. (a) illustrates the general framework of most time series models. (b) illustrates two core modules modified when applying CCM to base models: cluster assigner and cluster-aware Feed Forward. The cluster assigner learns channel clustering based on intrinsic similarities and creates prototype embeddings for each cluster via a cross-attention mechanism. The clustering probabilities $\{p_{i,k}\}$ are subsequently used in the cluster-aware Feed Forward to average $\{\mathbf{\theta}_{k}\}_{k=1}^{K}$, which are layer weights assigned to $K$ clusters, obtaining weights $\mathbf{\theta}^{i}$ for the $i$-th channel. The learned prototypes retain pre-trained knowledge, enabling zero-shot forecasting on unseen samples in both univariate and multivariate scenarios.
Time series forecasting has been attracting a surge of interest across diverse fields, ranging from economics, energy (Alvarez et al., 2010; Koprinska et al., 2018), weather (Angryk et al., 2020; Lam et al., 2023), to transportation planning (Li et al., 2015; Yin & Shang, 2016). The complexity of the task is heightened by factors including seasonality, trend, noise in the data, and potential cross-channel information.
Despite the numerous deep learning time series models proposed recently (Chen et al., 2023; Zeng et al., 2022; Wen et al., 2022; Zhou et al., 2022b, 2021; Wu et al., 2021, 2022; Liu et al., 2022a), an unresolved challenge persists in the effective management of channel interaction within the forecasting framework (Li et al., 2023a; Peiwen & Changsheng, 2023). Previous works have explored two primary channel strategies: Channel-Independent (CI) and Channel-Dependent (CD) strategies. Channel-Independent (CI) strategy has shown promise in better forecasting performance by having individual models for each channel. However, a critical drawback is its limited generalizability and robustness on unseen channels (Han et al., 2023). Besides, it tends to overlook potential interactions between various channels. Conversely, Channel-Dependent (CD) strategy models all channels as a whole and captures intricate channel relations, while they tend to show oversmoothing and have trouble fitting to individual channels, especially when the similarity between channels is very low. Moreover, existing models typically treat univariate data in a CI manner, neglecting the interconnections between time series samples, even though these dependencies are commonly observed in real-world scenarios, such as stock market or weather forecasting (Ni & Xu, 2023; Yin et al., 2021; Jang & Matteson, 2023).
Proposed work. To address the aforementioned challenges, we propose a Channel Clustering Module (CCM) that balances individual channel treatment and captures necessary cross-channel dependencies simultaneously. CCM is motivated by the following key observations: All models, CI and CD, heavily rely on channel identity information. The level of reliance is anti-correlated with the similarity between channels (see Sec. 4.1 for an analysis). This intriguing phenomenon alludes to the model’s analogous behavior on similar channels. The proposed CCM thereby involves the strategic clustering of channels into cohesive clusters, where intra-cluster channels exhibit a higher degree of similarity, effectively replacing channel identity with cluster identity. To capture the underlying time series patterns within these clusters, we employ cluster-aware Feed Forward to ensure individual management of each cluster. Moreover, CCM learns expressive prototype embeddings in the training phase, which enable zero-shot forecasting on unseen samples by grouping them into appropriate clusters.
CCM is a plug-and-play solution that is adaptable to most mainstream time series models. We evaluate the effectiveness of CCM on four different time series backbones (aka. base models), TSMixer (Chen et al., 2023), DLinear (Zeng et al., 2022), PatchTST (Nie et al., 2022), and TimesNet (Wu et al., 2022). Extensive experiments verify the superiority of CCM in long-term and short-term forecasting benchmarks, achieving an average margin of $2.4\%$ and $7.2\%$, respectively. Additionally, we collect stock data from a diverse range of companies to construct a new stock univariate dataset. Leveraging information from intra-cluster samples, CCM consistently shows a stronger ability to accurately forecast stock prices within the dynamic and intricate stock market. Moreover, CCM enhances zero-shot forecasting capacities of time series backbones in cross-domain scenarios, which further highlights the robustness and versatility of CCM.
The contributions of this paper are: (1) We propose a novel and unified channel strategy, i.e., CCM, which is adaptable to most mainstream time series models. CCM explores the optimal trade-off between channel individual treatment and cross-channel modeling, (2) Through learning prototypes from clusters, CCM enables zero-shot forecasting on unseen samples in both univariate and multivariate scenarios, and (3) CCM demonstrates superiority in improving performance on long-term and short-term forecasting.
2 Related Work
2.1 Time Series Forecasting Models
Traditional machine learning methods such as Prophet (Taylor & Letham, 2018; Triebe et al., 2021), ARIMA (Zhang, 2003) capture the trend component and seasonality in time series (Ahmed et al., 2010). As data availability continues to grow, deep learning methods revolutionized this field, introducing more complex and efficient models (Torres et al., 2021; Lim & Zohren, 2021). Convolutional Neural Networks (CNNs) (Wu et al., 2022; Liu et al., 2022a; Bai et al., 2018; Wan et al., 2019; Sen et al., 2019), have been widely adopted to capture local temporal dependencies. Recurrent Neural Networks (RNNs) (Hewamalage et al., 2021; Rangapuram et al., 2018; Smyl, 2020; Salinas et al., 2020; Bai et al., 2018) excel in capturing sequential information, yet they often struggle with longer sequences. Transformer-based models (Zhou et al., 2021; Liu et al., 2022b; Wu et al., 2021; Nie et al., 2022; Zhang & Yan, 2022; Liu et al., 2023b; Wen et al., 2022; Tang & Matteson, 2021; Zhou et al., 2022b; Liu et al., 2021; Feng et al., 2024), typically equipped with self-attention mechanisms (Vaswani et al., 2017), demonstrate their proficiency in handling long-range dependencies, although they require substantial computational resources. Recently, linear models (Oreshkin et al., 2019; Zhang et al., 2022; Das et al., 2023), e.g., DLinear (Zeng et al., 2022), TSMixer (Chen et al., 2023), have gained popularity for their simplicity and effectiveness in long-term time series forecasting, but they may underperform with non-linear and complex patterns. Besides, traditional tricks, including trend-seasonal decomposition (Zeng et al., 2022; Cleveland et al., 1990; Wen et al., 2020) and multi-periodicity analysis (Benaouda et al., 2006; Percival & Walden, 2000; Wu et al., 2022; Wang et al., 2022; Zhang et al., 2024; Yi et al., 2023; Zhou et al., 2022a) continue to play a crucial role in aiding in the preprocessing stage for advanced models.
2.2 Channel Strategies in Time Series Forecasting
Most deep learning models (Wu et al., 2021; Liu et al., 2021; Zhou et al., 2022b) adopt the Channel-Dependent (CD) strategy, aiming to harness the full spectrum of information across channels. Conversely, the Channel-Independent (CI) approaches (Nie et al., 2022; Zeng et al., 2022) build forecasting models for each channel independently. Prior works on CI and CD strategy (Han et al., 2023; Li et al., 2023a; Montero-Manso & Hyndman, 2021) present that CI has high capacity and low robustness, whereas CD is the opposite. Predicting residuals with regularization (PRReg) (Han et al., 2023) is thereby proposed to incorporate a regularization term in the objective to encourage smoothness in future forecasting. However, the essential challenge of channel strategies from the model design perspective has not been solved and it remains challenging to develop a balanced channel strategy with both high capacity and robustness.
Prior research has explored effective clustering of channels to improve the predictive capabilities in diverse applications, including image classification (Jiang et al., 2010), natural language processing (NLP) (Marin et al., 2023; George & Sumathy, 2023), anomaly detection (Li et al., 2012; Syarif et al., 2012; Gunupudi et al., 2017). For instance, in traffic prediction (Ji et al., 2023; Liu et al., 2023a), clustering techniques have been proposed to group related traffic regions to capture intricate spatial patterns. Despite the considerable progress in these areas, the potential and effect of channel clustering in time series forecasting remain under-explored.
3 Preliminaries
Time Series Forecasting. Formally, let $X=[\boldsymbol{x}_{1},\ldots\boldsymbol{x}_{T}]\in\mathbb{R}^{T\times C}$ be a time series, where $T$ is the length of historical data. $\boldsymbol{x}_{t}\in\mathbb{R}^{C}$ represents the observation at time $t$. $C$ denotes the number of variates (i.e., channels). In the univariate case (i.e., $C=1$), the series is composed of single observations at each time step, whereas in the multivariate case (i.e., $C>1$), each $\boldsymbol{x}_{t}$ is a vector of multiple observations. The objective is to construct a predictive model $f$ that estimates the future values of the series, $Y=[\hat{\boldsymbol{x}}_{T+1},\ldots,\hat{\boldsymbol{x}}_{T+H}]\in\mathbb{R}^{H\times C}$, where $H$ is the forecasting horizon. We use $X_{[:,i]}\in\mathbb{R}^{T}$ ($X_{i}$ for simplicity) to denote the $i$-th channel in the time series.
Channel Dependent (CD) and Channel Independent (CI). The CI strategy models each channel $X_{i}$ separately and ignores the potential channel interactions. This approach is typically denoted as $f^{(i)}:\mathbb{R}^{T}\rightarrow\mathbb{R}^{H}$ for $i=1,\cdots,C$, where $f^{(i)}$ gets access to the historical data specific to the $i$-th channel. In contrast, the CD strategy models all the channels as a whole, using a function $f:\mathbb{R}^{T\times C}\rightarrow\mathbb{R}^{H\times C}$. This strategy is essential in scenarios where channels are not just parallel data streams but are interrelated, such as in financial markets or traffic flows.
4 Proposed Method
In this work, we propose Channel Clustering Module (CCM), a model-agnostic method that is adaptable to most mainstream models. The pipeline of applying CCM is visualized in Figure 1. General time series models, shown in Figure 1(a), typically consist of three core components (Li et al., 2023a, b): an optional normalization layer (e.g., RevIN (Kim et al., 2021), SAN (Liu et al., 2023c)), temporal modules including linear layers, transformer-based, or convolutional backbones, and a feed-forward layer that forecasts the future values. Motivated by the empirical observation discussed in Sec. 4.1, CCM presents with a cluster assigner preceding the temporal modules, followed by a cluster-aware Feed Forward (Sec. 4.2). The cluster assigner implements channel clustering based on intrinsic similarities and employs a cross-attention mechanism to generate prototypes for each cluster, which stores the knowledge from the training set and endows the model with zero-shot forecasting capacities.
4.1 Motivation for Channel Similarity
To motivate our similarity-based clustering method, we provide a toy experiment to verify the following two assumptions: (1) Existing forecasting methods heavily rely on channel identity information. (2) This reliance clearly anti-correlates with channel similarity: for channels with high similarity, channel identity information is less important. Together, these two assumptions motivate us to design an approach that provides cluster identity instead of channel identity, combining the best of both worlds: high capacity and generalizability.
Table 1: Averaged performance gain from a channel identity information ($\Delta\mathcal{L}(\%)$) and Pearson Correlation Coefficients (PCC) between $\{\Delta\mathcal{L}_{ij}\}_{i,j}$ and $\{\textsc{Sim}(X_{i},X_{j})\}_{i,j}$. The values are averaged across all test samples.
Base Model TSMixer DLinear PatchTST TimesNet
Channel Strategy CD CI CI CD
ETTh1 $\Delta\mathcal{L}(\%)$ 2.67 1.10 11.30 18.90
PCC - 0.67 - 0.66 - 0.61 - 0.66
ETTm1 $\Delta\mathcal{L}(\%)$ 4.41 5.55 6.83 14.98
PCC - 0.68 - 0.67 - 0.68 - 0.67
Exchange $\Delta\mathcal{L}(\%)$ 16.43 19.34 27.98 24.57
PCC - 0.62 - 0.62 - 0.47 - 0.49
We select four recent and popular time series models with different backbones. TSMixer (Chen et al., 2023) and DLinear are linear models. PatchTST (Nie et al., 2022) is a transformer-based model with a patching mechanism and TimesNet (Wu et al., 2022) is a convolutional network that captures multi-periodicity in data. Among these, TSMixer and TimesNet utilize a Channel-Dependent strategy while DLinear and PatchTST adopt the Channel-Independent design. We train a time series model across all channels and evaluate the channel-wise Mean Squared Error (MSE) loss on the test set. Then, we repeat training while randomly shuffling channels in each batch. Note that for both CD and CI models, this means channel identity information will be removed. We report the average performance gain across all channels based on the random shuffling experiments (denoted as $\Delta\mathcal{L}(\%)$) in Table 1. We see that all models rely on channel identity information to achieve better performance.
Next, we show that this performance gap anti-correlates with channel similarity. We define channel similarity based on radial basis function kernels (Zhang et al., 2004) as
$\textsc{Sim}(X_{i},X_{j})=\text{exp}(\frac{-\|X_{i}-X_{j}\|^{2}}{2\sigma^{2}}),\vspace{-0.1cm}$ (1)
where $\sigma$ is a scaling factor. Note that the similarity is computed on the standardized time series to avoid scaling differences. More details are discussed in Appendix A. The performance difference in MSE from the random shuffling experiment for channel $i$ is denoted as $\Delta\mathcal{L}_{i}$. We define $\Delta\mathcal{L}_{ij}:=|\Delta\mathcal{L}_{i}-\Delta\mathcal{L}_{j}|$ and calculate the Pearson Correlation Coefficients between $\{\Delta\mathcal{L}_{ij}\}_{i,j}$ and $\{\textsc{Sim}(X_{i},X_{j})\}_{i,j}$, as shown in Table 1, consistently verifying our assumption.
4.2 CCM: Channel Clustering Module
Channel Clustering. Motivated by the above observations, we first initialize a set of $K$ cluster embedding $\{c_{1},\cdots,c_{K}\}$, where $c_{k}\in\mathbb{R}^{d}$, $d$ is the hidden dimension and $K$ is a hyperparameter. Given a multivariate time series $X\in\mathbb{R}^{T\times C}$, each channel in the input $X_{i}$ is transformed into a $d$-dimensional channel embedding $h_{i}$ through an MLP. The probability that a given channel $X_{i}$ is associated with the $k$-th cluster is the normalized inner-product of the cluster embedding $c_{k}$ and the channel embedding $h_{i}$, which is computed as
$p_{i,k}=\text{Normalize}(\frac{c_{k}^{\top}h_{i}}{\|c_{k}\|\|h_{i}\|})\in[0,1].\vspace{-0.1cm}$ (2)
The normalization operator ensures that $\sum_{k}p_{i,k}=1$ and validates the clustering probability distribution across $k$ clusters. We utilize reparameterization trick (Jang et al., 2016) to obtain the clustering membership matrix $\mathbf{M}\in\mathbb{R}^{C\times K}$ where $\mathbf{M}_{ik}\approx\text{Bernoulli}(p_{i,k})$. Higher probability $p_{i,k}$ results in $\mathbf{M}_{ik}$ close to 1, leading to deterministic existence of certain channel in the corresponding cluster.
Prototype Learning. The cluster assigner also creates a $d$-dimensional prototype embedding for each cluster in the training phase. Let $\mathbf{C}=[c_{1},\cdots,c_{K}]\in\mathbb{R}^{K\times d}$ denote the cluster embedding, and $\mathbf{H}=[h_{1},\cdots,h_{C}]\in\mathbb{R}^{C\times d}$ denote the hidden embedding of the channels. To emphasize the intra-cluster channels and remove interference from out-of-cluster channel information, we design a modified cross-attention as follows,
$\widehat{\mathbf{C}}=\text{Normalize}\left(\text{exp}(\frac{(W_{Q}\mathbf{C})(W_{K}\mathbf{H})^{\top}}{\sqrt{d}})\odot\mathbf{M}^{\top}\right)W_{V}\mathbf{H},$ (3)
where the clustering membership matrix $\mathbf{M}$ is an approximately binary matrix to enable sparse attention on intra-cluster channels specifically. $W_{Q}$, $W_{K}$ and $W_{V}$ are learnable parameters. The prototype embedding $\widehat{\mathbf{C}}\in\mathbb{R}^{K\times d}$ serves as the updated cluster embedding for subsequent clustering probability computing in Eq. 2.
Cluster Loss. We further introduce a specifically designed loss function for the clustering quality, termed ClusterLoss, which incorporates both the alignment of channels with their respective clusters and the distinctness between different clusters in a self-supervised learning context. Let $\mathbf{S}\in\mathbb{R}^{C\times C}$ denote the channel similarity matrix with $\mathbf{S}_{ij}=\textsc{Sim}(X_{i},X_{j})$ defined in Eq. 1. The ClusterLoss is formulated as:
$\mathcal{L}_{C}=-\operatorname{Tr}\left(\mathbf{M}^{\top}\mathbf{S}\mathbf{M}\right)+\operatorname{Tr}\left(\left(\mathbf{I}-\mathbf{M}\mathbf{M}^{\top}\right)\mathbf{S}\right),$ (4)
where $\operatorname{Tr}$ indicates a trace operator. $\operatorname{Tr}\left(\mathbf{M}^{\top}\mathbf{S}\mathbf{M}\right)$ maximizes the channel similarities within clusters, which is a fundamental requirement for effective clustering. $\operatorname{Tr}\left(\left(\mathbf{I}-\mathbf{M}\mathbf{M}^{\top}\right)\mathbf{S}\right)$ instead encourages separation between clusters, which further prevents overlap and ambiguity in clustering assignments. $\mathcal{L}_{C}$ captures meaningful time series prototypes without relying on external labels or annotations. The overall loss function thereby becomes $\mathcal{L}=\mathcal{L}_{F}+\beta\mathcal{L}_{C}$, where $\mathcal{L}_{F}$ is the general forecasting loss such as MSE loss; and $\beta$ is a regularization parameter for a balance between forecasting accuracy and cluster quality.
Cluster-aware Feed Forward. Instead of using individual Feed Forward per channel in a CI manner or sharing one Feed Forward across all channels in a CD manner, we assign a separate Feed Forward to each cluster to capture the underlying shared time series patterns within the clusters. In this way, we use cluster identity to replace channel identity. Each Feed Forward is parameterized with a single linear layer due to its efficacy in time series forecasting (Zeng et al., 2022; Li et al., 2023a; Chen et al., 2023). Let $h_{\theta_{k}}(\cdot)$ represent the linear layer for the $k$-th cluster with weights $\theta_{k}$. $Z_{i}$ represents the hidden embedding of the $i$-th channel before the last layer. The final forecast is thereby averaged across the outputs of all cluster-aware Feed Forward with $\{p_{i,k}\}$ as weights, e.g., $Y_{i}=\sum_{k}p_{i,k}h_{\theta_{k}}(Z_{i})$ for the $i$-th channel. For computational efficiency, it is equivalent to $Y_{i}=h_{\theta^{i}}(Z_{i})$ with averaged weights $\theta^{i}=\sum_{k}p_{i,k}\theta_{k}$.
Univariate Adaptation. In the context of univariate time series forecasting, we extend the proposed method to clustering on samples. We leverage the similarity between two univariate time series as defined in Eq. 1, and classify univariate time series with comparable patterns into the same cluster. This univariate adaptation allows it to capture interrelation within samples and extract valuable insights from analogous time series. This becomes particularly valuable in situations where meaningful dependencies exist among various univariate samples, such as the stock market.
Zero-shot Forecasting. The prototype embeddings acquired during the training phase serve as a compact representation of the pre-trained knowledge and can be harnessed for seamless knowledge transfer to unseen samples or new channels. We apply this knowledge to unseen instances by computing the clustering probability distribution on the pre-trained clusters, following Eq. 2. The clustering probability is subsequently used for averaging cluster-aware Feed Forward. We disable the cross-attention operation to fix the prototype embeddings in zero-shot forecasting scenarios. It is worth noting that zero-shot forecasting is adaptable to both univariate and multivariate scenarios. We refer to Appendix B for detailed discussion.
4.3 Complexity Analysis
CCM effectively strikes a balance between the CI and CD strategies. On originally CI models, CCM introduces strategic clustering on channels, which not only reduces the model complexity but also enhances their generalizability. Simultaneously, CCM increases the model complexity on originally CD models with negligible overhead for higher capacities. We refer to Figure 5 in Appendix C.5 for empirical analysis. Theoretically, the computational complexity of clustering probability computation (Eq. 2) and the cross-attention (Eq. 3) are $\mathcal{O}(KCd)$, where $K,C$ are the number of clusters and channels, respectively, and $d$ is the hidden dimension. One may also use more attention mechanisms (Kitaev et al., 2020; Wang et al., 2020; Shen et al., 2021) for efficiency. The complexity of cluster-aware Feed Forward scales linearly in $C,K$, and the forecasting horizon $H$.
5 Experiments
CCM consistently improves performance based on CI or CD models by significant margins across multiple benchmarks and settings, including long-term forecasting on 9 public multivariate datasets (Sec. 5.2); short-term forecasting on 2 univariate datasets (Sec. 5.3); and zero-shot forecasting in cross-domain and cross-granularity scenarios (Sec. 5.4).
5.1 Experimental Setup
Datasets. For long-term forecasting, we experiment on 9 popular benchmarking datasets across diverse domains (Zhou et al., 2021; Wu et al., 2021; Lai et al., 2018), including weather, traffic and electricity. M4 dataset (Makridakis et al., 2018) is used in short-term forecasting, which is a univariate dataset that covers time series across diverse domains and various sampling frequencies from hourly to yearly. We further provide a new stock time series dataset with 1390 univariate time series. Each time series records the price history of an individual stock spanning 10 years. Due to the potential significant fluctuations in stock performance across different companies, this dataset poses challenges for models capturing diverse and evolving stock patterns in financial markets. The statistics of long-term and short-term datasets are shown in Table 3 and Table 3.
Table 2: The statistics of datasets
in long-term forecasting. Horizon
is $\{96,192,336,720\}$.
Dataset Channels Length Frequency
ETTh1&ETTh2 7 17420 1 hour
ETTm1&ETTm2 7 69680 15 min
ILI 7 966 1 week
Exchange 8 7588 1 day
Weather 21 52696 10 min
Electricity 321 26304 1 hour
Traffic 862 17544 1 hour
Table 3: Dataset details of M4 and Stock in short-term forecasting.
Dataset Length Horizon
M4 Yearly 23000 6
M4 Quarterly 24000 8
M4 Monthly 48000 18
M4 Weekly 359 13
M4 Daily 4227 14
M4 Hourly 414 48
Stock (Ours) 10000 7/24
We follow standard protocols (Zhou et al., 2021; Wu et al., 2021, 2022) for data splitting on public benchmarking datasets. As for the stock dataset, we divide the set of stocks into train/validation/test sets with a ratio of 7:2:1. Therefore, validation/test sets present unseen samples (i.e., stocks) for model evaluation. This evaluation setting emphasizes the data efficiency aspect of time series models for scenarios where historical data is limited or insufficient for retraining from scratch given unseen instances. More details on datasets are provided in Appendix C.1.
Base Models and Experimental Details. CCM is a model-agnostic channel strategy that can be applied to arbitrary time series forecasting models for improved performance. We meticulously select four recent state-of-the-art time series models as base models: TSMixer (Chen et al., 2023), DLinear (Zeng et al., 2022), PatchTST (Nie et al., 2022) and TimesNet (Wu et al., 2022), which mainly cover three mainstream paradigms, including linear models, transformer-based and convolutional models. For fair evaluation, we use the optimal experiment configuration as provided in the official code to implement both base models and the enhanced version with CCM. All the experiments are implemented by PyTorch on a single NVIDIA RTX A6000 48GB GPU. Experiment configurations and implementations are detailed in Appendix C.3. Experimental results in the following sections are averaged on five runs with different random seeds. Refer to Appendix C.6 for standard deviation results.
Table 4: Long-term forecasting results on 9 real-world datasets in terms of MSE and MAE, the lower the better. The forecasting horizons are $\{96,192,336,720\}$. The better performance in each setting is shown in bold. The best results for each row are underlined. The last column shows the average percentage of MSE/MAE improvement of CCM over four base models.
Model TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM IMP(%)
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.361 0.392 0.365 0.393 0.375 0.399 0.371 0.393 0.375 0.398 0.371 0.396 0.384 0.402 0.380 0.400 0.539
192 0.404 0.418 0.402 0.418 0.405 0.416 0.404 0.415 0.415 0.425 0.414 0.424 0.436 0.429 0.431 0.425 0.442
336 0.422 0.430 0.423 0.430 0.445 0.440 0.438 0.443 0.422 0.440 0.417 0.429 0.491 0.469 0.485 0.461 0.908
720 0.463 0.472 0.462 0.470 0.489 0.488 0.479 0.497 0.449 0.468 0.447 0.469 0.521 0.500 0.520 0.493 0.333
ETTm1 96 0.285 0.339 0.283 0.337 0.299 0.343 0.298 0.343 0.294 0.351 0.289 0.338 0.338 0.375 0.335 0.371 1.123
192 0.339 0.365 0.336 0.368 0.335 0.365 0.334 0.365 0.334 0.370 0.333 0.363 0.374 0.387 0.373 0.383 0.482
336 0.361 0.406 0.359 0.393 0.370 0.386 0.365 0.385 0.373 0.397 0.370 0.392 0.410 0.411 0.412 0.416 0.716
720 0.445 0.470 0.424 0.421 0.427 0.423 0.424 0.417 0.416 0.420 0.419 0.430 0.478 0.450 0.477 0.448 1.852
ETTh2 96 0.284 0.343 0.278 0.338 0.289 0.353 0.285 0.348 0.278 0.340 0.274 0.336 0.340 0.374 0.336 0.371 1.371
192 0.339 0.385 0.325 0.393 0.384 0.418 0.376 0.413 0.341 0.382 0.339 0.355 0.402 0.414 0.400 0.410 1.806
336 0.361 0.406 0.361 0.399 0.442 0.459 0.438 0.455 0.329 0.384 0.327 0.383 0.452 0.452 0.449 0.445 0.823
720 0.445 0.470 0.438 0.464 0.601 0.549 0.499 0.496 0.381 0.424 0.378 0.415 0.462 0.468 0.457 0.461 4.370
ETTm2 96 0.171 0.260 0.167 0.260 0.167 0.260 0.166 0.258 0.174 0.261 0.168 0.256 0.187 0.267 0.189 0.270 0.860
192 0.221 0.296 0.220 0.296 0.284 0.352 0.243 0.323 0.238 0.307 0.231 0.300 0.249 0.309 0.250 0.310 3.453
336 0.276 0.329 0.277 0.330 0.369 0.427 0.295 0.358 0.293 0.346 0.275 0.331 0.321 0.351 0.318 0.347 6.012
720 0.420 0.422 0.369 0.391 0.554 0.522 0.451 0.456 0.373 0.401 0.374 0.400 0.408 0.403 0.394 0.391 7.139
Exchange 96 0.089 0.209 0.085 0.206 0.088 0.215 0.085 0.214 0.094 0.216 0.088 0.208 0.107 0.234 0.105 0.231 2.880
192 0.195 0.315 0.177 0.300 0.178 0.317 0.171 0.306 0.191 0.311 0.185 0.309 0.226 0.344 0.224 0.340 3.403
336 0.343 0.421 0.312 0.405 0.371 0.462 0.300 0.412 0.343 0.427 0.342 0.423 0.367 0.448 0.361 0.442 5.875
720 0.898 0.710 0.847 0.697 0.966 0.754 0.811 0.683 0.888 0.706 0.813 0.673 0.964 0.746 0.957 0.739 5.970
ILI 24 1.914 0.879 1.938 0.874 2.215 1.081 1.935 0.935 1.593 0.757 1.561 0.750 2.317 0.934 2.139 0.936 4.483
36 1.808 0.858 1.800 0.851 2.142 0.977 1.938 0.942 1.768 0.794 1.706 0.780 1.972 0.920 1.968 0.914 2.561
48 1.797 0.873 1.796 0.867 2.335 1.056 2.221 1.030 1.799 0.916 1.774 0.892 2.238 0.940 2.229 0.937 1.602
60 1.859 0.895 1.810 0.876 2.479 1.088 2.382 1.096 1.850 0.943 1.735 0.880 2.027 0.928 2.041 0.930 2.491
Weather 96 0.149 0.198 0.147 0.194 0.192 0.250 0.187 0.245 0.149 0.198 0.147 0.197 0.172 0.220 0.169 0.215 1.729
192 0.201 0.248 0.192 0.242 0.248 0.297 0.240 0.285 0.194 0.241 0.191 0.238 0.219 0.261 0.215 0.257 2.539
336 0.264 0.291 0.244 0.281 0.284 0.335 0.274 0.324 0.244 0.282 0.245 0.285 0.280 0.306 0.274 0.291 2.924
720 0.320 0.336 0.318 0.334 0.339 0.374 0.320 0.357 0.320 0.334 0.316 0.333 0.365 0.359 0.366 0.362 1.476
Electricity 96 0.142 0.237 0.139 0.235 0.153 0.239 0.142 0.247 0.138 0.233 0.136 0.231 0.168 0.272 0.158 0.259 2.480
192 0.154 0.248 0.147 0.246 0.158 0.251 0.152 0.248 0.153 0.247 0.153 0.248 0.184 0.289 0.172 0.262 3.226
336 0.163 0.264 0.161 0.262 0.170 0.269 0.168 0.267 0.170 0.263 0.168 0.262 0.198 0.300 0.181 0.284 2.423
720 0.208 0.300 0.204 0.299 0.233 0.342 0.230 0.338 0.206 0.296 0.210 0.301 0.220 0.320 0.205 0.309 1.417
Traffic 96 0.376 0.264 0.375 0.262 0.411 0.284 0.411 0.282 0.360 0.249 0.357 0.246 0.593 0.321 0.554 0.316 1.488
192 0.397 0.277 0.340 0.279 0.423 0.287 0.422 0.286 0.379 0.256 0.379 0.254 0.617 0.336 0.562 0.331 3.175
336 0.413 0.290 0.411 0.289 0.438 0.299 0.436 0.297 0.401 0.270 0.389 0.255 0.629 0.336 0.579 0.341 2.120
720 0.444 0.306 0.441 0.302 0.467 0.316 0.471 0.318 0.443 0.294 0.430 0.281 0.640 0.350 0.587 0.366 1.445
5.2 Long-term Forecasting Results
We report the mean squared error (MSE) and mean absolute error (MAE) on nine real-world datasets for long-term forecasting evaluation in Table 3. The forecasting horizon is $\{96,192,336,720\}$. From the table, we observe that the model enhanced with CCM outperforms the base model in general. Specifically, CCM improves the long-term forecasting performance in $90.27\%$ cases in MSE and $84.03\%$ cases in MAE across 144 different experiment settings. Remarkably, CCM achieves a substantial boost on DLinear, with a significant reduction on MSE by $5.12\%$ and MAE by $3.04\%$. The last column of the table quantifies the average percentage improvement in terms of MSE/MAE, which underscores the consistent enhancement brought by CCM across all forecasting horizons and datasets.
Comparison with Regularization Solution. PRReg (Han et al., 2023) recently displays to improve the performance of CD models, by predicting residuals with a regularization term in the training objective. We evaluate the effectiveness of PRReg and our proposed CCM on long-term forecasting performance enhancement based on CI and CD models. Following the previous training setting (Han et al., 2023), we develop CI and CD versions for Linear (Zeng et al., 2022) and Transformer (Vaswani et al., 2017) and report MSE loss. Table 5 shows that CCM surpasses PRReg in most cases, highlighting its efficacy compared with regularization solutions. See full results in Appendix C.4.
[…]
CD CI +PRReg +CCM
ETTh1 Linear 0.402 0.345 0.342 0.342
Transformer 0.861 0.655 0.539 0.518
ETTm1 Linear 0.404 0.354 0.311 0.310
Transformer 0.458 0.379 0.349 0.300
Weather Linear 0.142 0.169 0.131 0.130
Transformer 0.251 0.168 0.180 0.164
ILI Linear 2.343 2.847 2.299 2.279
Transformer 5.309 4.307 3.254 3.206
Electricity Linear 0.195 0.196 0.196 0.195
Transformer 0.250 0.185 0.185 0.183
5.3 Short-term Forecasting Results
Table 6: Short-term forecasting results on M4 dataset in terms of SMAPE, MASE, and OWA, the lower the better. The better performance in each setting is shown in bold.
Model TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM IMP(%)
Yearly SMAPE 14.702 14.676 16.965 14.337 13.477 13.304 15.378 14.426 7.286
MASE 3.343 3.370 4.283 3.144 3.019 2.997 3.554 3.448 9.589
OWA 0.875 0.873 1.058 0.834 0.792 0.781 0.918 0.802 11.346
Quaterly SMAPE 11.187 10.989 12.145 10.513 10.380 10.359 10.465 10.121 6.165
MASE 1.346 1.332 1.520 1.243 1.233 1.224 1.227 1.183 7.617
OWA 0.998 0.984 1.106 0.931 0.921 0.915 0.923 0.897 6.681
Monthly SMAPE 13.433 13.407 13.514 13.370 12.959 12.672 13.513 12.790 2.203
MASE 1.022 1.019 1.037 1.005 0.970 0.941 1.039 0.942 4.238
OWA 0.946 0.944 0.956 0.936 0.905 0.895 0.957 0.891 3.067
Others SMAPE 7.067 7.178 6.709 6.160 4.952 4.643 6.913 5.218 10.377
MASE 5.587 5.302 4.953 4.713 3.347 3.128 4.507 3.892 7.864
OWA 1.642 1.536 1.487 1.389 1.049 0.997 1.438 1.217 9.472
Avg. SMAPE 12.867 12.807 13.639 12.546 12.059 11.851 12.880 11.914 5.327
MASE 1.887 1.864 2.095 1.740 1.623 1.587 1.836 1.603 10.285
OWA 0.957 0.948 1.051 0.917 0.869 0.840 0.955 0.894 6.693
Table 7: Short-term forecasting results on stock dataset in terms of MAE and MSE, the lower the better. The forecasting horizon is $\{7,24\}$. The better performance in each setting is shown in bold.
Horizon Metric TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM IMP(%)
7 MSE 0.939 0.938 0.992 0.883 0.896 0.892 0.930 0.915 3.288
MAE 0.807 0.806 0.831 0.774 0.771 0.771 0.802 0.793 2.026
24 MSE 1.007 0.991 0.996 0.917 0.930 0.880 0.998 0.937 5.252
MAE 0.829 0.817 0.832 0.781 0.789 0.765 0.830 0.789 3.889
In both M4 and stock datasets, we follow the univariate forecasting setting. For M4 benchmarks, we adopt the evaluation setting in prior works (Oreshkin et al., 2019) and report the symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA) in Table 6. As for the stock dataset, we implement MAE and MSE as metrics in Table 7. See more details on metrics in Appendix C.2.
Remarkably, the efficacy of CCM is consistent across all M4 sub-datasets with different sampling frequencies. Specifically, CCM outperforms the state-of-the-art linear model (DLinear) by a significant margin of $11.62\%$, and outperforms the best convolutional method TimesNet by $8.88\%$. In table 7, we also observe a significant performance improvement on the stock dataset, achieved by applying CCM. In the stock dataset, we test on new samples (i.e., univariate stock time series) that the model has not seen in the training phase to evaluate the model’s generalization and robustness. By memorizing cluster-specific knowledge from analogous samples, the model potentially captures various market trends and behaviors and thereby makes more accurate and informed forecasting.
5.4 Zero-shot Forecasting Results
Table 8: Zero-shot forecasting results on ETT datasets. The forecasting horizon is $\{96,720\}$. The best value in each row is underlined.
Model TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM IMP(%)
Generalization Task MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
[…]
720 0.374 0.414 0.370 0.413 0.569 0.549 0.520 0.517 0.414 0.442 0.386 0.423 0.540 0.508 0.516 0.491 4.326
[…]
720 1.252 0.815 1.215 0.803 1.881 0.948 1.138 0.809 1.459 0.845 1.249 0.795 1.623 0.981 1.601 0.964 10.249
[…]
720 1.765 0.982 1.758 0.980 2.091 1.061 1.681 0.954 1.925 1.014 1.718 0.966 2.204 1.031 1.874 1.012 7.824
[…]
720 0.695 0.584 0.540 0.519 0.511 0.518 0.484 0.502 1.010 0.968 0.936 0.686 1.274 0.783 0.845 0.642 16.243
[…]
720 1.472 0.872 1.464 0.866 1.813 0.938 1.253 0.844 1.832 1.052 1.532 0.863 1.861 1.016 1.671 0.967 11.439
[…]
720 2.275 1.137 1.754 1.065 1.992 1.001 1.740 0.970 2.651 1.149 1.695 0.971 2.183 1.103 1.742 0.983 15.952
Existing time series models tend to be rigidly tailored to specific domains, leading to poor generalization on data from unseen domains. In contrast, CCM leverages learned prototypes to capture cluster-specific knowledge. This enables meaningful comparisons between unseen time series and pre-trained knowledge, facilitating accurate zero-shot forecasting. Following prior work (Jin et al., 2023), we adopt ETT collections (Zhou et al., 2021), where ETTh1 and ETTh2 are hourly recorded, while ETTm1 and ETTm2 are minutely recorded. “1” and “2” indicate two different regions where the datasets originated.
Table 8 shows MSE and MAE results on test datasets. CCM consistently improves the zero-shot forecasting capacity of base time series models in 48 scenarios, including generalization to different domains and different granularities. Specifically, based on the results, we make the following observations. (1) CCM exhibits more significant performance improvement with longer forecasting horizons, highlighting the efficacy of memorizing and leveraging pre-trained knowledge in zero-shot forecasting scenarios. (2) Achieving cross-granularity (i.e., different sampling frequencies) generalization presents greater challenges than cross-domain (i.e., different sampling regions) generalization. However, the result comparison between task 111 and 222 (or task 444 and 555 ) shows that CCM exhibits superiority in more challenging cross-granularity generalization scenarios. (3) CCM demonstrates a better effect on originally CI base models. For instance, the averaged improvement rates on two CI models, i.e., DLinear and PatchTST, are $10.48\%$ and $11.13\%$ respectively, while the improvement rates on TSMixer and TimesNet are $5.14\%$ and $9.63\%$. Overall, the experimental results verify the superiority and efficacy of CCM in enhancing zero-shot transfer forecasting.
5.5 Qualitative Visualization
Channel Clustering Visualization. Figure 2 presents the t-SNE visualization of channel and prototype embeddings within ETTh1 and ETTh2 datasets with DLinear as the base model. Each point represents a channel within a sample, with varying colors indicating different channels. In ETTh1 dataset, we discern a pronounced clustering of channels 0, 2, and 4, suggesting that they may be capturing related or redundant information within the dataset. Concurrently, channels 1, 3, 5, and 6 coalesce into another cluster. The similarity matrix in the lower left further corroborates these findings. Clustering is also observable in ETTh2 dataset, particularly among channels 0, 4, and 5, as well as channels 2, 3, and 6. Comparatively, channel 1 shows a dispersion among clusters, partly due to its capturing of unique or diverse aspects of the data that do not closely align with the features represented by any clusters. The clustering results demonstrate that CCM not only elucidates the intricate relationships and potential redundancies among the channels but also offers critical insights for feature analysis and enhancing the interpretability of time series models.
Figure 2: t-SNE visualization of channel and prototype embedding given by DLinear with CCM on (a) ETTh1 and (b) ETTh2 dataset. The lower left corner shows the similarity matrix between channels.
Weight Visualization of Cluster-aware Projection. Figure 3 depicts the weights visualization for the cluster-aware Feed Forward on ETTh1 and ETTm1 datasets, revealing distinct patterns that are indicative of the model’s learned features (Li et al., 2023a; Zeng et al., 2022; Yi et al., 2023). For instance, in the ETTm1 dataset, Cluster 0 shows bright diagonal striping patterns, which may suggest that it is primarily responsible for capturing the most dominant periodic signals in the corresponding cluster. In contrast, Cluster 1 exhibits denser stripes, indicating its role in refining the representation by capturing more subtle or complex periodicities that the first layer does not. The visualization implies the model’s ability to identify and represent periodicity in diverse patterns, which is crucial for time-series forecasting tasks that are characterized by intricate cyclic behaviors.
(a) ETTh1 Dataset
(b) ETTm1 Dataset
Figure 3: Weights visualization of cluster-wise linear layers on (a) ETTh1 and (b) ETTm1 datasets. The input and output lengths are 336 and 96, respectively. We observe the different periodicities captured by different clusters.
5.6 Ablation Study on Cluster Number
Figure 4: Ablation Study on Cluster Ratios in terms of MSE loss with four base models. The forecasting horizon is 96. (left: ETTh1 dataset; right: ETTm1 dataset)
Figure 4 shows an ablation study on cluster ratios, which is defined as the ratio of the number of clusters to the number of channels. $0.0$ means all channels are in a single cluster. We observe that the MSE loss slightly decreases and then increases as the cluster ratio increases, especially for DLinear, PatchTST, and TimesNet. Time series models with CCM achieve the best performance when the cluster ratio is in the range of $[0.2,0.6]$. It is worth noticing that DLinear and PatchTST, two CI models among four base models, benefit consistently from channel clustering with any number of clusters. More results are provided in Appendix D.
6 Conclusion
This work introduces a novel Channel Clustering Module (CCM) to address the challenge of effective channel management in time series forecasting. CCM strikes a balance between individual channel treatment and capturing cross-channel dependencies by clustering channels based on their intrinsic similarity. Extensive experiments demonstrate the efficacy of CCM in multiple benchmarks. It would be valuable to investigate the applicability of CCM in other domains beyond time series forecasting in future works.
Impact Statement. This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Channel Similarity
Essentially, the similarity between two time series $X_{i}$ and $X_{j}$ is defined as $\textsc{SIM}(X_{i},X_{j})=\text{exp}(\frac{-d(X_{i},X_{j})}{2\sigma^{2}})$, where $d(\cdot,\cdot)$ can be any distance metric (Chen & Ng, 2004; Ding et al., 2008), such as Euclidean Distance ($L_{p}$), Editing Distance (ED) and Dynamic Time Warping (DTW) (Gold & Sharir, 2018). One may also use other similarity definitions, such as Longest Common Subsequence (LCSS) and Cross-correlation (CCor).
Firstly, the choice of Euclidean distance in this work is motivated by its efficiency and low computational complexity, especially in the case of large datasets or real-time applications. Let $H$ denote the length of the time series. The complexity of the above similarity computation is shown in Table 9.
Table 9: Complexity of similarity computation
Euclidean Edit Distance DTW LCSS CCor
$\mathcal{O}(H)$ $\mathcal{O}(H^{2})$ $\mathcal{O}(H^{2})$ $\mathcal{O}(H^{2})$ $\mathcal{O}(H^{2})$
Secondly, it’s worth noting that while there are various similarity computation approaches, studies have demonstrated a strong correlation between Euclidean distance and other distance metrics (Kljun & Tersˇek, 2020). This high correlation suggests that, despite different mathematical formulations, these metrics often yield similar results when assessing the similarity between time series. This empirical evidence supports the choice of Euclidean distance as a reasonable approximation of similarity for practical purposes. In our implementation, we select $\sigma=5$ in Eq. 1 to compute the similarities based on Euclidean distance.
Appendix B Multivariate and Univariate Adaptation
We provide pseudocodes for training time series models enhanced with CCM in Algorithm 1. Algorithm 2 displays pseudocodes for the inference phase, where both the training and test sets have the same number of channels. The components in the pretrained model $\mathcal{F}$, highlighted in blue, remain fixed during the inference phase. It’s important to note that zero-shot forecasting in Algorithm 2 is adaptable to various scenarios. Let’s discuss these scenarios:
  • Training on a univariate dataset and inferring on either univariate or multivariate samples: In this case, the model learns prototypes from a vast collection of univariate time series in the training set. As a result, the model can effortlessly adapt to forecasting unseen univariate time series in a zero-shot manner. To forecast unseen multivariate time series, we decompose each multivariate sample into multiple univariate samples, where each univariate sample can be processed by the pretrained model. The future multivariate time series can be obtained by stacking multiple future univariate time series.
  • Training on a multivariate dataset and inferring on either univariate or multivariate samples: For Channel-Dependent models, test samples should have the same number of channels as the training samples, as seen in sub-datasets within ETT collections (Zhou et al., 2021). In contrast, for Channel-Independent models, zero-shot forecasting can be performed on either univariate or multivariate samples, even when they have different numbers of channels.
Algorithm 1 Forward function of time series models with channel clustering module. $C$ is the number of channels in the dataset. $K$ is the number of clusters. $T$ is the length of historical data. $H$ is the forecasting horizon.
  Input: Historical time series $X\in\mathbb{R}^{T\times C}$
  Output: Future time series $Y\in\mathbb{R}^{H\times C}$; Prototype embedding $\mathbf{C}\in\mathbb{R}^{K\times d}$
  Initialize the weights of $K$ linear layer $\theta_{k}$ for $k=1,\cdots,K$
  Initialize $K$ cluster embedding $c_{k}\in\mathbb{R}^{d}$ for $k=1,\cdots,K$$\triangleright$ Cluster Embedding $\mathbf{C}$
  $X\leftarrow\text{Normalize}(X)$
  $\mathbf{S}_{i,j}\leftarrow\text{exp}(\frac{-\|X_{i}-X_{j}\|^{2}}{2\sigma^{2}})$. $\triangleright$ Compute Similarity Matrix $\mathbf{S}$
  $h_{i}\leftarrow\text{MLP}(X_{i})$ for each channel $i$. $\triangleright$ Channel Embedding $\mathbf{H}$ via MLP Encoder in the Cluster Assigner
  $p_{i,k}\leftarrow\text{Normalize}(\frac{c_{k}^{\top}h_{i}}{\|c_{k}\|\|h_{i}\|})\in[0,1].$$\triangleright$ Compute Clustering Probability Matrix $\mathbf{P}$
  $\mathbf{M}\leftarrow\text{Bernoulli}(\mathbf{P})$$\triangleright$ Sample Clustering Membership Matrix $\mathbf{M}$
  $\mathbf{C}\leftarrow\text{Normalize}\left(\text{exp}(\frac{(W_{Q}\mathbf{C})(W_{K}\mathbf{H})^{\top}}{\sqrt{d}})\odot\mathbf{M}^{\top}\right)W_{V}\mathbf{H}$. $\triangleright$ Update Cluster Embedding $\mathbf{C}$ via Cross Attention
[…]
  for channel $i$ in $\{1,2,\cdots,C\}$ do
     $Y_{i}\leftarrow h_{\theta^{i}}(\hat{\mathbf{H}}_{i})$ where $\theta^{i}=\sum_{k}p_{i,k}\theta_{k}$. $\triangleright$ Weight Averaging and Projection
  end for
Algorithm 2 Zero-shot forecasting of time series models with channel clustering module. $C$ is the number of channels in both the training and test datasets. $K$ is the number of clusters. $T$ is the length of historical data. $H$ is the forecasting horizon.
  Input: Historical time series $X\in\mathbb{R}^{T\times C}$; Pretrained Models $\mathcal{F}$
  Output: Future time series $Y\in\mathbb{R}^{H\times C}$;
  Load the weights of $K$ linear layer $\theta_{k}$ for $k=1,\cdots,K$ from $\mathcal{F}$
  Load $K$ cluster embedding $c_{k}$$\in\mathbb{R}^{d}$ for $k=1,\cdots,K$ from $\mathcal{F}$$\triangleright$ Cluster Embedding $\mathbf{C}$
  $X\leftarrow\text{Normalize}(X)$
  $\mathbf{S}_{i,j}\leftarrow\text{exp}(\frac{-\|X_{i}-X_{j}\|^{2}}{2\sigma^{2}})$. $\triangleright$ Compute Similarity Matrix $\mathbf{S}$
[…]
  $p_{i,k}\leftarrow\text{Normalize}(\frac{c_{k}^{\top}h_{i}}{\|c_{k}\|\|h_{i}\|})\in[0,1].$$\triangleright$ Compute Clustering Probability Matrix $\mathbf{P}$
  $\mathbf{M}\leftarrow\text{Bernoulli}(\mathbf{P})$$\triangleright$ Sample Clustering Membership Matrix $\mathbf{M}$
[…]
  for channel $i$ in $\{1,2,\cdots,C\}$ do
     $Y_{i}\leftarrow h_{\theta^{i}}(\hat{\mathbf{H}}_{i})$ where $\theta^{i}=\sum_{k}p_{i,k}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\theta_{k}}$. $\triangleright$ Weight Averaging and Projection
  end for
Appendix C Experiments
C.1 Datasets
Public Datasets. We utilize nine commonly used datasets for long-term forecasting evaluation. Firstly, ETT collection (Zhou et al., 2021), which documents the oil temperature and load features of electricity transformers over the period spanning July 2016 to July 2018. This dataset is further subdivided into four sub-datasets, ETThs and ETTms, with hourly and 15-minute sampling frequencies, respectively. s can be $1$ or $2$, indicating two different regions. Electricity dataset (Trindade, 2015) encompasses electricity consumption data from 321 clients from July 2016 to July 2019. Exchange (Lai et al., 2017) compiles daily exchange rate information from 1990 to 2016. Traffic dataset contains hourly traffic load data from 862 sensors in San Francisco areas from 2015 to 2016. Weather dataset offers a valuable resource with 21 distinct weather indicators, including air temperature and humidity, collected every 10 minutes throughout the year 2021. ILI documents the weekly ratio of influenza-like illness patients relative to the total number of patients, spanning from 2002 to 2021. Dataset statistics can be found in Table 10.
We adopt M4 dataset for short-term forecasting evaluation, which involves 100,000 univariate time series collected from different domains, including finance, industry, etc.The M4 dataset is further divided into 6 sub-datasets, according to the sampling frequency.
Table 10: The statistics of dataset in long-term and short-term forecasting tasks
Tasks Dataset Channels Forecast Horizon Length Frequency Domain
Long-term ETTh1 7 $\{96,192,336,720\}$ 17420 1 hour Temperature
ETTh2 7 $\{96,192,336,720\}$ 17420 1 hour Temperature
ETTm1 7 $\{96,192,336,720\}$ 69680 15 min Temperature
ETTm2 7 $\{96,192,336,720\}$ 69680 15 min Temperature
Illness 7 $\{96,192,336,720\}$ 966 1 week Illness Ratio
Exchange 8 $\{96,192,336,720\}$ 7588 1 day Exchange Rates
Weather 21 $\{96,192,336,720\}$ 52696 10 min Weather
Electricity 321 $\{96,192,336,720\}$ 26304 1 hour Electricity
Traffic 862 $\{96,192,336,720\}$ 17544 1 hour Traffic Load
Short-term M4-Yearly 1 6 23000 yearly Demographic
M4 Quarterly 1 8 24000 quarterly Finance
M4 Monthly 1 18 48000 monthly Industry
M4 Weekly 1 13 359 weekly Macro
M4 Daily 1 14 4227 daily Micro
M4 Hourly 1 48 414 hourly Other
Stock 1 $\{7,24\}$ 10000 5 min Stock
Stock Dataset. We design a new time series benchmark dataset constructed from publicly available stock-market data. We deploy commercial stock market API to probe the market statistics for 1390 stocks spanning 10 years from Nov.25, 2013 to Sep.1, 2023. We collect stock price data from 9:30 a.m. to 4:00 p.m. every stock open day except early closure days. The sampling granularity is set to be 5 minutes. Missing record is recovered by interpolation from nearby timestamps and all stock time series are processed to have aligned timestep sequences. We implement short-term forecasting on the stock dataset, which is more focused on market sentiment, and short-term events that can cause stock prices to fluctuate over days, weeks, or months. Thereby, we set the forecasting horizon as $7$ and $24$.
C.2 Metrics
Following standard evaluation protocols (Wu et al., 2022), we utilize the Mean Absolute Error (MAE) and Mean Square Error(MSE) for long-term and stock price forecasting. The Symmetric Mean Absolute Percentage Error (SMAPE), Mean Absolute Scaled Error (MASE), and Overall Weighted Average (OWA) are used as metrics for M4 dataset (Makridakis et al., 2018; Oreshkin et al., 2019). The formulations are shown in Eq. 5. Let $\mathbf{y}_{t}$ and $\widehat{\mathbf{y}}_{t}$ denote the ground-truth and the forecast at the $t$-th timestep, respectively. $H$ is the forecasting horizon. In M4 dataset, MASE is a standard scale-free metric, where $s$ is the periodicity of the data (e.g., 12 for monthly recorded sub-dataset) (Makridakis et al., 2018). MASE measures the scaled error w.r.t. the naïve predictor that simply copies the historical records of $s$ periods in the past. Instead, SMAPE scales the error by the average between the forecast and ground truth. Particularly, OWA is an M4-specific metric (Team et al., 2018) that assigns different weights to each metric.
$\text{MAE}=\frac{1}{H}\sum_{t=1}^{H}\left|\mathbf{y}_{t}-\widehat{\mathbf{y}}_{t}\right|,\qquad\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(\mathbf{y}_{t}-\widehat{\mathbf{y}}_{t})^{2},$ (5)
$\text{SMAPE}=\frac{200}{H}\sum_{i=1}^{H}\frac{\left|\mathbf{y}_{t}-\widehat{\mathbf{y}}_{t}\right|}{\left|\mathbf{y}_{t}\right|+\left|\widehat{\mathbf{y}}_{t}\right|},\qquad\text{MASE}=\frac{1}{H}\sum_{i=1}^{H}\frac{\left|\mathbf{y}_{t}-\widehat{\mathbf{y}}_{t}\right|}{\frac{1}{H-s}\sum_{j=s+1}^{H}\left|\mathbf{y}_{j}-\mathbf{y}_{j-s}\right|},$
OWA=12[ $\text{OWA}=\frac{1}{2}\left[\frac{\text{ SMAPE }}{\text{ SMAPE }_{\text{Naïve2 }}}+\frac{\text{ MASE }}{\text{MASE}_{\text{Naïve2 }}}\right]$
C.3 Experiment Details
To verify the superiority of CCM in enhancing the performance of mainstream time series models, we select four popular and state-of-the-art models, including linear models such as TSMixer (Chen et al., 2023), DLinear (Zeng et al., 2022), transformer-based model PatchTST (Nie et al., 2022) and convolution-based model TimesNet (Wu et al., 2022). We build time series models using their official codes and optimal model configuration [^1] https://github.com/yuqinie98/PatchTST [^2] https://github.com/cure-lab/LTSF-Linear [^3] https://github.com/google-research/google-research/tree/master/tsmixer [^4] https://github.com/thuml/TimesNet.
In the data preprocessing stage, we apply reversible instance normalization (Kim et al., 2021) to ensure zero mean and unit standard deviation, avoiding the time series distribution shift. Forecasting loss is MSE for long-term forecasting datasets and the stock dataset. Instead, we use SMAPE loss for M4 dataset. We select Adam (Kingma & Ba, 2014) with the default hyperparameter configuration for $(\beta_{1},\beta_{2})$ as (0.9, 0.999). An early-stopping strategy is used to mitigate overfitting. The experiments are conducted on a single NVIDIA RTX A6000 48GB GPU, with PyTorch (Paszke et al., 2019) framework. The code is available at the following anonymous link: https://anonymous.4open.science/r/TimeSeriesCCM-ECC2. We use the official codes and follow the best model configuration to implement the base models. Then we apply CCM to the base models, keeping the hyperparameters unchanged for model backbones. Experiment configurations, including the number of MLP layers in the cluster assigner, the layer number in the temporal module, hidden dimension, the best cluster number, and regularization parameter $\beta$ on nine real-world datasets are shown in Table 11.
Table 11: Experiment configuration.
# clusters $\beta$ linear layers in MLP hidden dimension # layers (TSMixer) # layers (PatchTST) # layers (TimesNet)
ETTh1 2 0.3 1 128 2 2 3
ETTm1 2 0.3 1 64 2 4 2
ETTh2 2 0.3 1 64 2 4 3
ETTm2 2 0.9 1 24 2 4 4
Exchange 2 0.9 1 32 2 4 3
ILI 2 0.9 1 36 2 6 3
Weather [2,5] 0.5 2 64 4 3 3
Electricity [3,10] 0.5 2 128 4 3 3
Traffic [3,10] 0.5 2 128 4 3 3
C.4 Comparison between CCM and Other Approach
Table 12: Full Results on Comparison between CCM and existing regularization method for enhanced performance on CI/CD strategies in terms of MSE metric. The best results are highlighted in bold.
CD CI +PRReg +CCM
ETTh1(48) Linear 0.402 0.345 0.342 0.342
Transformer 0.861 0.655 0.539 0.518
ETTh2(48) Linear 0.711 0.226 0.239 0.237
Transformer 1.031 0.274 0.273 0.284
ETTm1(48) Linear 0.404 0.354 0.311 0.310
Transformer 0.458 0.379 0.349 0.300
ETTm2(48) Linear 0.161 0.147 0.136 0.146
Transformer 0.281 0.148 0.144 0.143
Exchange(48) Linear 0.119 0.051 0.042 0.042
Transformer 0.511 0.101 0.044 0.048
Weather(48) Linear 0.142 0.169 0.131 0.130
Transformer 0.251 0.168 0.180 0.164
ILI(24) Linear 2.343 2.847 2.299 2.279
Transformer 5.309 4.307 3.254 3.206
Electricity(48) Linear 0.195 0.196 0.196 0.195
Transformer 0.250 0.185 0.185 0.183
Predict Residuals with Regularization. Prior work (Han et al., 2023) demonstrates that the main drawback of CD models is their inclination to generate sharp and non-robust forecasts, deviating from the actual trend. Thereby, Predict Residuals with Regularization (PRReg for simplicity), a specifically designed regularization objective, is proposed to improve the robustness of CD methods as follows.
$\mathcal{L}=\frac{1}{N}\sum_{j=1}^{N}\mathcal{L}_{F}\left(f\left(X^{(j)}-A^{(j)}\right)+A^{(j)},Y^{(j)}\right)+\lambda\Omega(f),$ (6)
where the superscript $j$ indicates the sample index. $\mathcal{L}_{F}$ is MSE loss. $A^{(j)}=X^{j}_{L,:}$ represents the last values of each channel in $X^{(j)}$. Therefore, the objective encourages the model to generate predictions that are close to the nearby historical data and keep the forecasts smooth and robust. The regularization term $\Omega$ ($L_{2}$ norm in practice) further restricts the complexity of the model and ensures smoothness in the predictions. It was demonstrated that PRReg can achieve better performance than original CD and CI strategies (Han et al., 2023). We conduct extensive experiments on long-term forecasting benchmarks to compare PRReg and CCM. The full results are shown in Table 12. The numbers in parentheses next to the method represent the forecasting horizon. We set the length of the look-back window to 36 for ILI and 96 for other datasets for consistency with previous works (Han et al., 2023). The base models are linear model (Zeng et al., 2022) and basic transformer (Vaswani et al., 2017). The values in the PRReg column are the best results with any $\lambda$, collected from (Han et al., 2023). We observe from Table 12 that CCM successfully improves forecasting performance on original CI/CD strategies (or reached comparable results) in 13 out of 16 settings, compared with PRReg method.
(a) Model Size Analysis on ETTh1 dataset
(b) Runtime Efficiency Analysis on ETTh1 dataset
Figure 5: Efficiency analysis in model size and running time on ETTh1 dataset.
C.5 Complexity and Efficiency Analysis
We evaluate the model size and runtime efficiency of the proposed CCM with various numbers of clusters on ETTh1 dataset, as shown in Figure 5. The batch size is 32, and the hidden dimension is 64. We keep all other hyperparameters consistent to ensure fair evaluation. It is worth noting that CCM reduces the model complexity based on Channel-Independent models (e.g., PatchTST, DLinear), since CCM essentially uses cluster identity to replace channel identity. The generalizability of CI models is thereby enhanced as well. When it comes to Channel-Dependent models, CCM increases the model complexity with negligible overhead, considering the improved forecasting performance.
C.6 Error Bar
Experimental results in this paper are averaged from five runs with different random seeds. We report the standard deviation for base models and CCM-enhanced versions on long-term forecasting datasets in Table 13, M4 dataset in Table 14 and stock dataset in Table 15.
Table 13: Standard deviation of Table 3 on long-term forecasting benchmarks. The forecasting horizon is 96.
ETTh1 ETTm1 ETTh2 ETTm2 Exchange ILI Weather Electricity Traffic
MSE 0.361$\pm$0.001 0.285$\pm$0.001 0.284$\pm$0.001 0.171$\pm$0.001 0.089$\pm$0.004 1.914$\pm$0.031 0.149$\pm$0.008 0.142$\pm$0.002 0.376$\pm$0.006
TSMixer MAE 0.392$\pm$0.001 0.339$\pm$0.001 0.343$\pm$0.002 0.260$\pm$0.001 0.209$\pm$0.009 0.879$\pm$0.009 0.198$\pm$0.009 0.237$\pm$0.004 0.264$\pm$0.005
+CCM MSE 0.365$\pm$0.001 0.283$\pm$0.002 0.278$\pm$0.001 0.167$\pm$0.001 0.085$\pm$0.006 1.938$\pm$0.015 0.147$\pm$0.007 0.139$\pm$0.005 0.375$\pm$0.006
MAE 0.393$\pm$0.001 0.337$\pm$0.002 0.338$\pm$0.002 0.260$\pm$0.001 0.206$\pm$0.011 0.874$\pm$0.012 0.194$\pm$0.009 0.235$\pm$0.008 0.262$\pm$0.005
MSE 0.375$\pm$0.002 0.299$\pm$0.001 0.289$\pm$0.001 0.167$\pm$0.001 0.088$\pm$0.006 2.215$\pm$0.031 0.192$\pm$0.011 0.153$\pm$0.004 0.411$\pm$0.006
DLinear MAE 0.399$\pm$0.001 0.343$\pm$0.001 0.353$\pm$0.001 0.260$\pm$0.001 0.215$\pm$0.010 1.081$\pm$0.009 0.250$\pm$0.008 0.239$\pm$0.005 0.284$\pm$0.005
+CCM MSE 0.371$\pm$0.001 0.298$\pm$0.001 0.285$\pm$0.001 0.166$\pm$0.002 0.085$\pm$0.006 1.935$\pm$0.034 0.187$\pm$0.015 0.142$\pm$0.003 0.411$\pm$0.005
MAE 0.393$\pm$0.001 0.343$\pm$0.002 0.348$\pm$0.02 0.258$\pm$0.002 0.214$\pm$0.013 0.935$\pm$0.012 0.245$\pm$0.020 0.247$\pm$0.006 0.282$\pm$0.004
MSE 0.375$\pm$0.003 0.294$\pm$0.003 0.278$\pm$0.003 0.174$\pm$0.003 0.094$\pm$0.008 1.593$\pm$0.016 0.149$\pm$0.008 0.138$\pm$0.004 0.360$\pm$0.005
PatchTST MAE 0.398$\pm$0.004 0.351$\pm$0.004 0.340$\pm$0.004 0.261$\pm$0.003 0.216$\pm$0.012 0.757$\pm$0.015 0.198$\pm$0.012 0.233$\pm$0.005 0.249$\pm$0.005
+CCM MSE 0.371$\pm$0.002 0.289$\pm$0.005 0.274$\pm$0.006 0.168$\pm$0.003 0.088$\pm$0.006 1.561$\pm$0.021 0.147$\pm$0.008 0.136$\pm$0.002 0.357$\pm$0.007
MAE 0.396$\pm$0.003 0.338$\pm$0.005 0.336$\pm$0.006 0.256$\pm$0.003 0.208$\pm$0.009 0.750$\pm$0.009 0.197$\pm$0.013 0.231$\pm$0.006 0.246$\pm$0.006
MSE 0.384$\pm$0.005 0.338$\pm$0.006 0.340$\pm$0.005 0.187$\pm$0.005 0.107$\pm$0.009 2.317$\pm$0.024 0.172$\pm$0.011 0.168$\pm$0.002 0.593$\pm$0.010
TimesNet MAE 0.402$\pm$0.005 0.375$\pm$0.006 0.374$\pm$0.005 0.267$\pm$0.003 0.234$\pm$0.013 0.934$\pm$0.010 0.220$\pm$0.013 0.272$\pm$0.005 0.321$\pm$0.008
+CCM MSE 0.380$\pm$0.004 0.335$\pm$0.005 0.336$\pm$0.003 0.189$\pm$0.003 0.105$\pm$0.006 2.139$\pm$0.038 0.169$\pm$0.015 0.158$\pm$0.003 0.554$\pm$0.009
MAE 0.400$\pm$0.004 0.371$\pm$0.006 0.371$\pm$0.005 0.270$\pm$0.005 0.231$\pm$0.010 0.936$\pm$0.018 0.215$\pm$0.024 0.259$\pm$0.006 0.316$\pm$0.008
Table 14: Standard deviation of Table 6 on M4 dataset
Model TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM
Yearly SMAPE 0.122 0.130 0.087 0.089 0.135 0.134 0.168 0.162
MASE 0.022 0.022 0.019 0.017 0.018 0.021 0.0017 0.017
OWA 0.002 0.002 0.002 0.002 0.007 0.009 0.010 0.011
Quaterly SMAPE 0.101 0.103 0.100 0.100 0.079 0.073 0.106 0.105
MASE 0.016 0.016 0.015 0.015 0.008 0.009 0.013 0.011
OWA 0.008 0.007 0.006 0.008 0.013 0.016 0.009 0.009
Monthly SMAPE 0.113 0.113 0.110 0.111 0.122 0.120 0.120 0.134
MASE 0.013 0.015 0.009 0.013 0.017 0.019 0.011 0.012
OWA 0.001 0.001 0.002 0.001 0.003 0.002 0.004 0.004
Others SMAPE 0.113 0.110 0.126 0.128 0.137 0.130 0.129 0.125
MASE 0.024 0.026 0.036 0.031 0.023 0.025 0.028 0.025
OWA 0.011 0.013 0.009 0.008 0.023 0.019 0.024 0.020
Avg. SMAPE 0.113 0.115 0.111 0.103 0.136 0.134 0.148 0.153
MASE 0.027 0.025 0.021 0.017 0.026 0.021 0.027 0.042
OWA 0.006 0.004 0.004 0.004 0.013 0.016 0.021 0.036
Table 15: Standard deviation of Table  7 on stock dataset
Horizon Metric TSMixer + CCM DLinear + CCM PatchTST + CCM TimesNet + CCM
7 MSE 0.001 0.001 0.001 0.001 0.003 0.003 0.004 0.004
MAE 0.001 0.001 0.001 0.001 0.003 0.002 0.003 0.003
24 MSE 0.002 0.002 0.001 0.002 0.005 0.004 0.005 0.005
MAE 0.001 0.001 0.001 0.001 0.003 0.002 0.003 0.004
Appendix D Ablation Study
D.1 Influence of Cluster Ratio
Table 16: Sensitivity of cluster ratio in terms of MSE metric. The forecasting horizon is 96. 0.0 means grouping all channels into the same cluster. original means the base model without the CCM mechanism.
Cluster Ratio Original 0.0 0.2 0.4 0.6 0.8 1.0
ETTh1 TSMixer 0.361 0.361 0.362 0.365 0.363 0.364 0.366
DLinear 0.375 0.378 0.371 0.371 0.372 0.372 0.371
PatchTST 0.375 0.380 0.372 0.371 0.373 0.376 0.375
TimesNet 0.384 0.384 0.380 0.383 0.385 0.385 0.388
ETTm1 TSMixer 0.285 0.285 0.283 0.283 0.284 0.285 0.286
DLinear 0.299 0.303 0.298 0.298 0.299 0.300 0.300
PatchTST 0.294 0.298 0.292 0.289 0.289 0.293 0.293
TimesNet 0.338 0.338 0.337 0.335 0.336 0.335 0.337
Exchange TSMixer 0.089 0.089 0.086 0.087 0.088 0.090 0.092
DLinear 0.088 0.093 0.088 0.087 0.085 0.089 0.089
PatchTST 0.094 0.095 0.089 0.088 0.088 0.091 0.093
TimesNet 0.107 0.107 0.105 0.105 0.107 0.107 0.107
Electricity TSMixer 0.142 0.142 0.139 0.139 0.140 0.143 0.143
DLinear 0.153 0.160 0.143 0.142 0.143 0.147 0.150
PatchTST 0.138 0.142 0.136 0.136 0.138 0.140 0.140
TimesNet 0.168 0.168 0.160 0.159 0.167 0.168 0.169
The number of clusters is an important hyperparameter in the CCM method. To verify the effectiveness of our design, we conduct an ablation study for all base models on four long-term forecasting datasets. The full results are shown in Table 16. We tune different cluster ratios, defined as the ratio of the number of clusters to the number of channels. Original means the original base model without any channel clustering mechanism. 0.0 indicates grouping all channels into the same cluster. We make the following observations. (1) For most cases, the channel clustering module (CCM) with any number of clusters greater than 1 consistently improves the forecasting performance upon base models. (2) For Channel-Independent base models, such as DLinear and PatchTST, grouping all channels into one cluster results in a channel fusion at the last layer, leading to a degradation in forecasting performance than the original CI models. (3) For most cases, the cluster ratio in the range of $[0.2,0.6]$ typically results in the best performance.
D.2 Influence of Look-back Window Length
Table 17: Short-term forecasting on stock dataset with different look-back window length in $\{14,21,28\}$. The forecasting length is 7. The best results with the same base model are underlined. Bold means CCM successfully enhances forecasting performance over the base model.
Forecast 7 24
Input 14 21 28 48 72 96
Length MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
TSMixer 0.947 0.806 0.974 0.816 0.939 0.807 1.007 0.829 1.016 0.834 1.100 0.856
+ CCM 0.896 0.778 0.954 0.808 0.938 0.806 0.991 0.817 1.006 0.824 1.078 0.851
DLinear 1.003 0.834 0.995 0.833 0.992 0.831 0.998 0.832 0.996 0.832 0.998 0.832
+ CCM 0.897 0.778 0.904 0.782 0.883 0.774 0.921 0.786 0.917 0.781 0.969 0.798
PatchTST 0.933 0.804 0.896 0.771 0.926 0.794 0.976 0.793 0.951 0.790 0.930 0.789
+ CCM 0.931 0.758 0.892 0.771 0.924 0.790 0.873 0.767 0.860 0.759 0.880 0.765
TimesNet 0.943 0.816 0.934 0.803 0.930 0.802 0.998 0.830 1.003 0.818 1.013 0.821
+ CCM 0.926 0.796 0.911 0.789 0.915 0.793 0.937 0.789 0.974 0.803 0.979 0.804
Imp. (%) 4.492 4.590 3.527 2.211 3.230 2.152 6.492 3.798 5.344 3.271 3.409 2.445
In this section, we conduct additional ablation experiments to investigate the effect of different look-back window lengths in the newly collected stock dataset, which determines how much historical information the time series model uses to make short-term forecasts. Specifically, the ablation study helps identify the risk of overfitting or underfitting based on the chosen look-back window length. An overly long window may lead to overfitting, while a short window may cause underfitting. Table 17 display the forecasting performance on forecasting horizon 7 and 24. The length of the look-back window ranges from two to four times the forecasting horizon. From Table 17, we make the following observations. (1) CCM effectively improves the base models’ forecasting performance in 24 cases across different base models, forecasting horizons, and look-back window lengths consistently. (2) Especially, CCM achieves better enhancement when the look-back window is shorter.
Appendix E Visualization Results
E.1 Channel-wise Performance and Channel Similarity
Figure 6 illustrates the channel-wise forecasting performance in terms of MSE metric and channel similarity on ETTh1 dataset with DLinear. We use the model’s performance difference on the original dataset and the randomly shuffled dataset as the model’s fitting ability on a specific channel. Note that MSE loss is computed on channels that have been standardized, which means that any scaling differences between them have been accounted for. Figure 6 highlights a noteworthy pattern: when two channels exhibit a higher degree of similarity, there tends to be a corresponding similarity in the performance on these channels. This observation suggests that channels with similar characteristics tend to benefit similarly from the optimization. It implies a certain level of consistency in the improvement of MSE loss when dealing with similar channels.
Figure 6: (a) Channel-wise forecasting performance and (b) Channel similarity on ETTh1 dataset illustrate the correlation between model performance and intrinsic similarity
Appendix F Discussion
This paper presents the Channel Clustering Module (CCM) for enhanced performance of time series forecasting models, aiming to balance the treatment of individual channels while capturing essential cross-channel interactions. Despite its promising contributions, there still exist limitations and directions for future work that warrant consideration.
Limitation. While CCM shows improvements in forecasting, its scalability to extremely large datasets remains to be tested. Moreover, the clustering and embedding processes in CCM introduce additional computational overhead. The efficiency of CCM in real-time forecasting scenarios, where computational resources are limited, requires further optimization.
Future Works. Future research can focus on adapting CCM to specific domains, such as biomedical signal processing or geospatial data analysis, where external domain-specific knowledge can be involved in the similarity computation. Furthermore, exploring alternative approaches to develop a dynamical clustering mechanism within CCM could potentially enhance forecasting efficacy. It is also worth examining the effectiveness of CCM in contexts beyond time series forecasting.